20.0-mL sample of 0.123 M diprotic acid (H2A) solution is titrated with 0.1013 M KOH. The acid ionization constants for the acid are Ka1=5.2×10−5 and Ka2=3.4×10−10.

Part A:

At what added volume of base does the first equivalence point occur?

Part B:

At what added volume of base does the second equivalence point occur?
Answer #1

Solution : - Given that - diprotic acid. (H2A) - 20.0ml of Sample kott = o-103 M Kai = 5.2x10-5 of 0.123 M Kaz - 3. H XLO ex1

> So what were the Pka and PKa2 values for??

miguelangelseguragiraldo Sat, Nov 13, 2021 11:34 AM